OpenStudy (anonymous):

At how many points on the interval -2pi to 2pi does the tangent to the graph of the curve y=xcosx have slope pi/2 i got the derivative to xcosx as xsinx+cosx... now do i set that equal to pi/2? if so how do you solve that because i dunno

OpenStudy (anonymous):

the derivative is cosx - xsinx

OpenStudy (anonymous):

thanks you! but what do i do after that?

OpenStudy (anonymous):

first note that f(x)=f(-x), for f(x)=cosx - xsinx consider only the positive x values, and multiply the final answer by 2 there is guaranteed to be at least one positive x solution, since at f(3pi/2)=3pi/2 > pi/2 and f(0)=1 and the function is continuous

OpenStudy (anonymous):

*for the first part, it should say f(pi)=-1 instead of f(0)=1

OpenStudy (anonymous):

the actual values of x are not easily solvable; the number of values is

OpenStudy (anonymous):

in summary f(0)=1 f(pi/2)=-pi/2 f(pi)=-1 f(3pi/2)=3pi/2 f(2pi)=1 the only domains we care about are (pi,3pi/2) and (3pi/2,2pi). f in these domain will hit pi/2 to verify that f only hits pi/2 once in each domain, we can compute f'(0)=0 f'(pi/2)=-2 f'(pi)=pi f'(3pi/2)=2 f'(5pi/3)=sqrt(3)-5pi/6 f'(2pi)=-2pi from pi to 3pi/2, f' is positive thus f is increasing from 5pi/3 to 2pi, f' is negative thus f is decreasing thus the number of points is 4